OpenStudy (anonymous):

What is the volume of a cylinder with the following dimensions:

OpenStudy (anonymous):

2pi*r^2*h, h=height

OpenStudy (anonymous):

that just confused me ._.

OpenStudy (anonymous):

are you familiar with the formula for the area of a circle?

OpenStudy (anonymous):

\[2\pi(r^2)\]

OpenStudy (anonymous):

i don't understand that .

OpenStudy (anonymous):

\[V= \pi*(r)^2*l\]

OpenStudy (anonymous):

r radius, l - length

OpenStudy (anonymous):

then read about it, http://www.mathsisfun.com/geometry/circle-area.html

OpenStudy (anonymous):

math has been really confusing for me .

OpenStudy (anonymous):

if you go to https://www.khanacademy.org/ make an account and go through their math program i guarantee math will start making sense, but it takes time and serious effort

OpenStudy (anonymous):

okay, i'll try it . thank you .

OpenStudy (anonymous):

as a rule of thumb for myself, every time i think "this is f****g stupid" it means im about to understand something.

OpenStudy (anonymous):

lololol you're funny . thanks for your help and the site to check out . greatly appreciate it .

OpenStudy (anonymous):

no problem, feel free to ask me any questions in the future

OpenStudy (anonymous):

okay , i sure will . (:

OpenStudy (anonymous):

okay , i've got a question ,

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

what do they mean by this ? Convert the mixed number, 1 3/4 , to a fraction. Reduce fraction if necessary.

OpenStudy (anonymous):

so think of 1 3/4 as "one whole and 3/4 of a whole" one whole in terms of 4 parts is 4/4 so you have 4/4 parts and 3/4 parts making..... total parts

OpenStudy (anonymous):

how would i put the problem into my phone , using my calculator on my phone ...

OpenStudy (anonymous):

3+4

OpenStudy (anonymous):

over 4

OpenStudy (anonymous):

thats easy .... i feel stupid.

OpenStudy (anonymous):

it's 4/7

OpenStudy (anonymous):

7/4

OpenStudy (anonymous):

OR 7/4

OpenStudy (anonymous):

you want to conceptualize it then you can solve any type of this question they throw at you. think of a fraction as parts of a whole, the upper number (called the numerator) is how many parts you have, and the lower number (called the denominator) is how many parts you need to make a whole. so 1 3/4 is one whole (4/4) and 3/4 of a whole making 3+7 parts out of a whole consisting of 4 parts or 7/4

OpenStudy (anonymous):

so doesn't that mean that it's a whole fraction ?

OpenStudy (anonymous):

yes, you have more than a whole. Theres usually, almost always, some sort of picture you can use to understand math concepts. For this one i think of pizzas, you basically have a whole pizza and three forths of a pizza

OpenStudy (anonymous):

oooohhhh that makes sense to me now ... but just a little . i understand what you're saying i'll just need to practice a little more to fully understand the concept of this.

OpenStudy (anonymous):

im sure youll do just fine

OpenStudy (anonymous):

i just have no clue how to work my next problem , it's about pi . i haven't really understood it .

OpenStudy (anonymous):

Pi is just a number that involves how many times you could wrap the diameter of a circle around the circle its self.

OpenStudy (anonymous):

ohh . could you explain more about it to me ?

OpenStudy (anonymous):



OpenStudy (anonymous):

i couldnt think of anything better than twizzlers as a unit of lenght, but you can use anything, a circle with a diameter of one jellybean, meter, toe nail, skittle, will have a circumfrence of pi jellybeans, meters, or skittles

OpenStudy (anonymous):

in math terms this is stated as "The ratio (fraction involving) the diameter and the circumference is equal to pi (3.14159)"

OpenStudy (anonymous):

or \[\frac{ C }{ D } = \pi =3.14159....., c= circumfrence, d= diameter\]